Phosphorous pentachloride decomposes to phosphorus trichloride according to this equation:



PCl5(l) <===> PCl(g) + Cl2(g). At equilibrium, [PC5] = 1.00 M and [Cl2] = 3.16 x 10^-2 M. Show your work.




A. Write the expression for determining the concentration of PCl3.





B. What is the equilibrium concentration of PCl3? Use: Keq = 1.00 x 10^-3.

1
Expert's answer
2022-05-04T13:38:02-0400

"PCl_5 \rightarrow PCl_3 + Cl_2"

"K_{eq}=\frac{[PCl_3][Cl_2]}{[PCl_5]}"

"[PCl_3]=\frac{K_{eq}[PCl_5]}{[Cl_2]}=\frac{1x10^{-3}x1}{3.16x10^{-2}}=3.16 M"